OpenStudy (anonymous):

Find the integral of:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{e ^{\frac{ 1 }{x }}dx }{ x^{3} }\]

OpenStudy (anonymous):

Let u= 1/x, du=1/x^2 ?

OpenStudy (anonymous):

Substitute \(1/x = u\) .

OpenStudy (anonymous):

\( du = -1/x^2 dx\)

OpenStudy (anonymous):

What is the form of integral in terms of u now ?

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ u du }{ x }\]

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

With the negative sign.. :)

OpenStudy (anonymous):

It should completely be in terms of u. There should be no x. But this is wrong !

OpenStudy (anonymous):

\[-\int\limits_{}^{}e^{u}du\]

OpenStudy (anonymous):

It should be \(-\int ue^u du\) .

hartnn (hartnn):

just to share an alternative method, you can also put \(\Large u = e^{\dfrac{1}{x}}\) but even then you will have to use product rule. so you can go ahead with your substitution :)

OpenStudy (anonymous):

Now apply integration by parts.

OpenStudy (anonymous):

\[-\frac{ u ^{2} }{ 2 } \left( e ^{u} \right)\]

OpenStudy (anonymous):

+c

OpenStudy (anonymous):

Not sure...

OpenStudy (anonymous):

The answer is wrong. Do know integration by parts ?

OpenStudy (anonymous):

Nope.. :)

OpenStudy (anonymous):

OK. If you have an integral like \(\int u(x)v(x) dx\), you can write it as \(\int U(x)V(x) dx = U(x) \int V(x) dx - \int \left(\frac{dU(x)}{dx} \int V(x)dx \right) dx\) Here x is u, U(u) is also u and V(u) = \(e^u\).

OpenStudy (anonymous):

Can you apply this ?

OpenStudy (anonymous):

Nope.. :( We haven't discuss that yet..

OpenStudy (anonymous):

Hmm. Try to work it out.

OpenStudy (anonymous):

So, \(\int u e^u du = u\int e^u du - \int \left(\frac{du}{du} \int e^udu\right)du = ue^u - \int e^udu = e^u(u-1)\) So \(- \int u e^u du = e^u(1-u)\) + const.